USE THE FOLLOWING SCENARIO FOR QUESTIONS 6-10. A 25-ml sample of 0.10 Macetic acid is titrated with 15-ml 0.10 M NaOH. (K, fo Can anyone check my answers, and correct me if I’m wrong?
Answer #1

6.

At equivalence point

mmoles of CH3COOH =  mmoles of NaOH

At 15 mL NaOH , the solution contains excess acetic acid.

hence equilbrium is due to dissociation of acetic acid.

CH3COOH(aq) + H2O(l)  \rightleftharpoons H3O+ (aq) + CH3COO- (aq)

Option A is correct.

7.

When moles of 15 ml 0.1 M NaOH is added

mmoles of NaOH = volume* molarity = 15*0.1= 1.5

mmoles acetic acid (AcOH) = (25*0.1) = 2.5

Excess mmoles of acid = (2.5 - 1.5) = 1

mmoles of sodium acetate (NaOAc) formed = moles of NaOH added = 1.5

Ka = 1.8*10-5, pKa = -log Ka = - log(1.8*10-5) = 4.74

Now the solution is an acid buffer

Using Hendersson -Hasselbalch equation.

pH = pKa + log [NaOAc/AcOH]

= 4.74 + log (1.5/1)

= 4.74 + 0.18 = 4.92 .

Correct answer is B.

8.

At half equivalence point

[NaOAc] = [AcOH]

pH = pKa

The point A is the half equivalence point.

9.

At the equivalence point , pH is due to dissociation of sodium acetate.

Hence equilibrium reaction is

CH3COO- (aq)  + H2O (l) \rightleftharpoons CH3COOH (aq) + OH- (aq).

10.

At equivalence point , pH of the solution is due to dissocistion of the salt (sodium acetate).

For salt of weak acid strong base , pH at equivalence point

pH = 7 + \frac{1}{2} (pka + logC)

at , equivalence point mmoles of salt = 25*0.1= 2.5

Volune of solution = (25+25)= 50.

Then , concentration of salt (c) = (2.5/50)= 0.05

Now,pH = 7 + \frac{1}{2} [4.74 + log(0.05)]

Or, pH = 7 + 1.72 = 8.72 .

Option C is correct.